# A Full AGN Feedback Prescription for Numerical Models: Negative, Positive and Hot Gas-Ejection Modes

### v1

@@ 3 Results
Before delving into the details of the analysis—namely, the genuine predictions of FEGA25 and comparisons with its predecessor and observational data—it is important to highlight a key point. In the analysis presented in C24, which examined various galaxy properties and scaling relations, FEGA24 successfully provided a robust description. However, FEGA25 not only achieves similar results but also introduces notable improvements, which will be the focus of the following discussion.
@@
Figure 2 presents the star formation rate (SFR)–stellar mass relation as predicted by FEGA24 (left panel) and FEGA25 (right panel), both compared to the observed main sequence of galaxies reported by Elbaz et al. (2007). Overall, both models successfully reproduce the main sequence and also provide a reasonable depiction of the broad red cloud. However, there is a significant difference between them. Specifically, while the main sequence predicted by FEGA24 is slightly biased high relative to the observed one—albeit still within the observational scatter—the main sequence predicted by FEGA25 aligns perfectly with the observed average (the contours represent different percentages of the data, as indicated by the colors in the accompanying bars, and the last contour encloses 90% of the full sample, while the remaining 10% are represented by individual black dots). This improvement is a direct outcome of the new AGN feedback implementation in FEGA25, which more effectively regulates the availability of hot gas compared to the mechanism employed in FEGA24.
@@
Another significant improvement over the previous model is illustrated in Figure 3, which shows the fraction of red galaxies as a function of their stellar mass. The predictions from FEGA25 are shown in blue, while those from FEGA24 are shown in green. Both are compared with observed fractions derived from SDSS-DR7 (De Lucia et al. 2024, represented by magenta and blue diamonds), SDSS-DR8 (Hirschmann et al. 2014, indicated by orange stars), and the GAMA survey (Davies et al. 2019, marked by purple triangles). Following C24, red galaxies are defined consistently with the above-mentioned observations, based on different thresholds in the logarithm of the specific star formation rate (SSFR): $-10.5$ for GAMA data, and $-10.6$ and $-11.0$ for SDSS data. The green and blue regions represent the range encompassing the highest and lowest cuts in $\log\text{SSFR}$.
@@
Figure 3 shows that FEGA25 reproduces the red fraction of galaxies in greater detail than FEGA24. This is evident not only in the overall trend of an increasing red fraction with stellar mass but particularly in its ability to match the observed fractions at low stellar masses. Interestingly, the drop in the red fraction at low stellar masses—although counterintuitive—is a direct consequence of the introduction of the hot gas ejection mode. Without this mechanism, FEGA25 would behave similarly to FEGA24, despite the differences in their positive feedback implementations. In low-mass halos, the ejection of gas appears to result in a larger fraction of dwarf galaxies remaining blue, likely due to the increased availability of cold gas. We will revisit this point in Section 4, where we will provide a comprehensive interpretation of the results.
However, a caveat must be noted. While FEGA24 generally falls between the various observational datasets across most of the stellar mass range, FEGA25 aligns very well with GAMA and SDSS-DR7 data, accurately capturing both trends and fractions, but deviates significantly from the SDSS-DR8 observations.
@@
The most evident difference between the two models lies in their trends with stellar mass. In FEGA24, the relation is flat, with a larger scatter for lower-mass galaxies and an average contribution of approximately 5%. In contrast, FEGA25 exhibits an increasing percentage and scatter with stellar mass, starting with virtually no contribution for dwarf and intermediate-mass galaxies and rising to about 5% for massive galaxies at $z=0$. At $z=3$, the percentages in FEGA24 remain consistent, albeit with larger scatter, while FEGA25 shows an increasing contribution toward lower-mass galaxies (with respect to $z=0$). This highlights a key distinction between the models: in FEGA24, redshift dependence is reflected only through the scatter, whereas in FEGA25, it is a clear and systematic trend.
@@
This result is particularly significant because it aligns well with recent theoretical predictions (e.g., Silk et al. 2024), which suggest that positive AGN feedback is stronger at high redshifts—when negative feedback mechanisms are less active—and diminishes at lower redshifts, as negative feedback becomes dominant. Nevertheless, given the limited observational evidence for positive feedback, especially at high redshifts, it is important to emphasize that neither implementation can be definitively ruled out. Both positive and negative feedback modes may coexist. The advantage of the FEGA25 prescription is its ability to accurately describe the redshift dependence of positive AGN feedback.
@@
To gain an understanding of the contribution of positive AGN feedback to star formation in FEGA25, we present the SFR histories of eight galaxies with increasing stellar mass, from top to bottom panels, in Figure 5. The red lines represent the total SFR histories, while the blue lines show the contribution from positive AGN feedback. As expected from the results in the previous figure, the overall contribution of the positive feedback mode to star formation increases with the galaxy’s stellar mass, both in terms of the mass it provides and the duration for which the positive feedback is active.
In lower-mass galaxies (top four panels), the contribution of positive AGN feedback is generally shorter and characterized by episodic bursts. In contrast, for higher-mass galaxies (bottom four panels), the SFR history driven by positive AGN feedback closely follows the total SFR. In all cases, which are representative examples, the positive feedback mode appears to be most efficient at high redshift and begins to fade as the negative feedback mode takes over cooling processes, reaching its maximum strength. The specific redshift at which this transition occurs varies from case to case. However, with the exception of a few occasional bursts, the positive mode begins to lose its effectiveness after $z=2$ in all cases. Another interesting feature visible in the bottom panels for the four massive galaxies is that the SFR due to positive feedback can reach, and in some instances exceed, 100 $M_{\odot}/\rm{yr}$.
@@
Finally, Figure 6 serves as the key figure of this study, clearly illustrating the role of our hot gas ejection mode. The plot shows, for each baryonic component (denoted by different colors as described in the legend), the fraction of mass relative to the total mass within the virial radius, normalized by the universal baryon fraction, as a function of halo mass, $\rm{Y}_{200}$. Solid lines represent the predictions of FEGA24, while dashed lines correspond to those of FEGA25. The hot gas fraction predicted by our models is compared with recent results from Popesso et al. (2024), shown as magenta squares, and a set of observational data from different authors (see caption), represented by dark blue circles. It is important to note that, in the latter case, the fraction was computed within the virial region using an overdensity of $\Delta=500$ (with halo masses converted to $M_{200}$ in the plot), rather than $\Delta=200$ as in our models, so these values should be considered as lower limits compared to ours  [^1] The relation between the hot gas fractions within the two overdensities can be approximated, in the absence of the density profile of the gas, by $\rm{Y}_{500}\sim\rm{Y}_{200}/1.15$ (e.g., Vikhlinin et al. 2006; Pratt et al. 2009; Popesso et al. 2024)..
@@
The key, and very important, distinction lies in the hot gas fraction (red lines), and by extension, in the ejected gas fraction (green lines). FEGA24 (solid line), which lacks a hot gas ejection mode in its AGN feedback, predicts a hot gas fraction of approximately 60%-65% for halos with $\log M_{\rm{halo}}>12.2$, from small groups to large clusters. In contrast, FEGA25 (dashed line), thanks to its hot gas ejection mode driven by AGN feedback, shows a considerably different hot gas fraction, more closely aligned with the observed data. The fraction of hot gas is noticeably reduced from low-mass to large groups, and while it matches the observations from Popesso et al. (2024) at the edge of the scatter in several data points, it represents a clear improvement over FEGA24.
Since the cold gas and stellar components remain largely unchanged, the excess gas is directed into the ejecta (green lines). This behavior aligns with the desired outcome, as highlighted by other authors (see the discussion in Popesso et al. 2024) and as discussed in Section 1. The hot gas ejection mode in the AGN feedback has proven to be effective, and in Section 4, we will explore the broader implications of our findings.
4 Discussion
The analysis presented above highlights several key features of the updated model FEGA25, made possible by the implementation of both the new positive AGN feedback and, more importantly, the novel hot gas ejection mode. In the following, we discuss each aspect of the improvements in detail, comparing them with observational evidence. Given the importance of the topics addressed, it is worth reiterating the fundamental differences between FEGA25 and its predecessor, FEGA24.
The first significant update in FEGA25 involves a more detailed description of stellar halos surrounding central galaxies (Contini et al. 2024a). In this framework, these stellar halos form through the reincorporation of stars originally belonging to the diffuse light or intracluster light (ICL) in a transition region between the ICL and the central galaxy. However, this implementation does not influence the analysis conducted in this work.
The second major update is particularly important and represents an unprecedented advancement in semianalytic modeling. While FEGA24 isolated the positive AGN feedback using Equation 4, FEGA25 introduces a more physically motivated approach, described by Equation 5. This new formulation links the positive AGN feedback to its negative counterpart, both controlled by a single free parameter, the BH accretion efficiency, $\kappa_{\rm{AGN}}$.
Both versions of our SAM successfully capture the primary properties of galaxies and many critical scaling relations investigated in C24. However, as demonstrated in the analysis above, FEGA25 introduces a range of improvements, bringing its predictions significantly closer to observational data. Setting aside the evolution of the SMF itself (as it was used during the model calibration), one key insight emerges from comparing the two models: the role and value of the parameter $\kappa_{\rm{AGN}}$.
In FEGA25, $\kappa_{\rm{AGN}}$ governs all three modes of AGN feedback, influencing BH growth, the BH response to cooling flows, the star formation bursts induced by cooling flows, and the amount of hot gas retained within the halo’s virial radius. The SMF evolution shows an intriguing result: although $\kappa_{\rm{AGN}}$ is 2.5 times lower in FEGA25, the model predicts reduced number densities at the high-mass end. This is unexpected since a lower $\kappa_{\rm{AGN}}$ would typically enhance the high-mass end of the SMF.
This result can be explained by the interconnection between the three feedback modes. The introduction of the third mode—hot gas ejection—ensures that some fraction of hot gas is expelled beyond the halo’s virial radius. While some of this gas is eventually reincorporated, the overall hot gas content within the halo decreases compared to a model without this mode (e.g., FEGA24). With a reduced amount of hot gas, cooling flows are naturally less massive, leading to less gas being accreted onto the BH. Consequently, $\kappa_{\rm{AGN}}$ can be lower in a model with hot gas ejection while still achieving the desired regulatory effects.
@@ 4 Discussion
The interconnection between the three feedback modes also has a remarkable and positive impact on the SFR of galaxies. As highlighted in the analysis above, FEGA25 provides a significantly improved description of the observed galaxy main sequence (Elbaz et al. 2007) and predicts red galaxy fractions more consistent with observations (Davies et al. 2019; De Lucia et al. 2024). Notably, this improvement in predicting the fraction of red galaxies is primarily driven by the ejection mode. As emphasized in Section 1 and reiterated here, this mechanism must effectively expel gas beyond the halo’s virial radius without disrupting the internal processes that govern the cold gas and stellar components, ensuring the preservation of galaxy properties already well described by the model. That said, we observed minor differences in the amount of cold gas and stars in low-mass halos, which notably enhance the model’s ability to capture the fraction of red dwarf galaxies.
A significant contribution to the overall success of FEGA25 also stems from the updated implementation of the positive feedback mode. Compared to the previous version, this new implementation accurately captures the expected redshift dependence: the mode is more effective at higher redshifts, where the negative mode remains largely inactive, and becomes weaker at lower redshifts as the negative mode increasingly suppresses cooling flows. The observed trend with stellar mass indicates that only intermediate- to high-mass galaxies experience a meaningful contribution from this mode throughout their evolution, though this contribution shifts toward lower mass scales at higher redshifts.
However, evaluating the reliability of the positive feedback mode is not straightforward due to the limited availability of observational data, particularly at high redshifts. Current predictions align with recent theoretical work (Silk et al. 2024 and references therein), but achieving a more comprehensive understanding will require additional observational data across a broader stellar mass range and higher redshifts.
@@
It is clear that the new achievements in our SAM stem from the introduction of a more comprehensive AGN feedback model. The idea that AGN can expel gas beyond the halo’s virial radius is novel in SAMs, though it has been quantified in numerical simulations (e.g., Pillepich et al. 2019; Angelinelli et al. 2022; Popesso et al. 2024, among the most recent studies). Popesso et al. (2024) provides an updated and detailed analysis of this phenomenon, comparing numerical simulation results with observational data (e.g., eROSITA eFEDS field data, Brunner et al. 2022). As discussed in Section 1, most simulations struggle to reproduce the observed fraction of hot gas within the virialized region over a broad range of halo masses, particularly at group scales. This shortfall underscores the need for an additional mechanism capable of pushing hot gas beyond the virial radius.
Our previous version, FEGA24, predicted hot gas fractions comparable to those found in the numerical simulations analyzed by Popesso et al. (2024), but these fractions do not match the observed trends. In contrast, the more advanced FEGA25 significantly reduces the hot gas fraction, especially for Milky Way-like halos and group-scale halos, where the mismatch between models and observations is most pronounced. This improvement highlights the importance of introducing a new AGN feedback mode in FEGA25, which expels hot gas beyond the halo’s virial radius, thereby narrowing the gap between predictions and observations. However, further investigation is needed for a particular reason. While FEGA25 performs considerably better than its predecessor, its predictions still tend toward the upper limits of the observed range, suggesting that an even more efficient mechanism may be required.
@@
As described in Section 2.1.3, we can assume that the residual energy, after reheating the cooling gas in the negative feedback mode, is sufficient to expel some of the hot gas beyond the virial radius. While this is a plausible and valuable approach, it represents an extreme scenario that warrants deeper exploration. We plan to address these questions in a forthcoming paper (Contini et al., in prep.), where we will focus on the evolution of the baryon content in halos and its dependence on time.

### v2

@@ 3.2 FEGA24 vs FEGA25
Before delving into the details of the analysis—namely, the genuine predictions of FEGA25 and comparisons with its predecessor and observational data—it is important to highlight a key point. In the analysis presented in C24, which examined various galaxy properties and scaling relations, FEGA24 showed a good level of reproduction. However, FEGA25 not only achieves similar results but also introduces notable improvements on some of them, which will be the focus of the following discussion.
@@
Figure 2 presents the star formation rate (SFR)–stellar mass relation at $z=0,1,2,3$, as predicted by FEGA24 (left panels) and FEGA25 (right panels), both compared to several observations as indicated by the legend. Overall, both models successfully reproduce the main sequence at $z=0$ and also provide a reasonable depiction of the broad red cloud. However, there is some difference between them. Specifically, while the main sequence predicted by FEGA24 is slightly biased high relative to the observed one—albeit still within the observational scatter—the main sequence predicted by FEGA25 aligns perfectly with the observed average (the contours represent different percentages of the data, as indicated by the colors in the accompanying bars, and the last contour encloses 90% of the full sample, while the remaining 10% are represented by individual black dots). This improvement is a direct outcome of the new AGN feedback implementation in FEGA25, which more effectively regulates the availability of hot gas compared to the mechanism employed in FEGA24.
At higher redshifts, the main sequence predicted by FEGA24 appears to be more consistent with observational data than that produced by FEGA25, which shows a closer alignment with other SAMs, such as L-Galaxies (Henriques et al. 2020, their Figure 7). However, it is worth noting that the main sequence from FEGA24 is generally broader compared to that of FEGA25. As we will discuss in more detail later, this broader distribution is a direct consequence of the stronger implementation of positive AGN feedback in FEGA24.
The overall evolution of the SFR is best represented by the cosmic SFR density as a function of redshift, shown in Figure 3. This provides a crucial benchmark for evaluating how well our models capture the buildup of stellar mass over cosmic time. In this figure, we compare the predictions of our models—depicted by the green line for FEGA24 and the blue line for FEGA25—against observational data from Driver et al. 2018 (red crosses) and D’Silva et al. 2023 (magenta diamonds).
A key result from this comparison is that FEGA24 successfully reproduces the cosmic SFR density only in the last $\sim$7 Gyr, particularly when compared to the dataset from D’Silva et al. (2023). However, at higher redshifts, its predictions deviate from the observational constraints, indicating that the original model struggles to fully capture the early phases of star formation history. In contrast, FEGA25 shows a much-improved agreement with the observed cosmic SFR density, particularly with the dataset from Driver et al. (2018). Notably, FEGA25 maintains a close match up to $z=5$, suggesting that the refinements in the new model allow for a more accurate representation of early galaxy formation and evolution.
@@
Another significant improvement over the previous model is illustrated in Figure 4, which shows the fraction of passive galaxies as a function of their stellar mass, and at different redshift (different panels). The predictions from FEGA25 are shown in blue, while those from FEGA24 are shown in green. Both are compared with observed fractions derived from Muzzin et al. 2013 (grey regions) and from Weaver et al. 2023 (red diamonds). Following C24, passive galaxies are defined consistently with the above-mentioned observations, based on different thresholds in the logarithm of the specific star formation rate (SSFR), set to be $0.3\cdot t^{-1}_{\rm{Hubble}}$ (around $10^{-11}{\rm yr}^{-1}$ at $z=0$). The green and blue regions represent the range encompassing the lowest and highest ($+0.5$ dex) thresholds in $\log\text{SSFR}$.
@@
Figure 4 shows both model can describe the evolution with redshift of the passive fraction of galaxies up to $z=2$, with FEGA25 being more accurate, but they both underestimate the observed fractions at $z=3$. FEGA25 reproduces the passive fraction of galaxies in greater detail than FEGA24 up to $z=2$. This is evident not only in the overall trend of an increasing passive fraction with stellar mass, but particularly in its ability to get closer to the observed fractions at lower stellar masses. Interestingly, the gap between the models in the passive fraction (at $z=0,1,2$) at low stellar masses—although counterintuitive—is a direct consequence of the introduction of the hot gas ejection mode. Without this mechanism, FEGA25 would behave similarly to FEGA24, despite the differences in their positive feedback implementations. In low-mass halos, the ejection of gas appears to result in a larger fraction of dwarf galaxies remaining active  [^3] This feature is caused by the introduction of the hot gas ejection mode. In fact, when FEGA25 is run with the new implementation of positive AGN feedback alone, the model still behaves similarly to the original version., likely due to the increased availability of cold gas (this point is under investigation in Contini et al. in prep).
Figure 5: As done in C24, we present the percentage of stellar mass contributed by the positive AGN feedback mode as a function of stellar mass (for central galaxies only). The predictions of FEGA24 (left panel) and FEGA25 (right panel) are shown at two redshifts: $z=0$ (green lines) and $z=3$ (black lines). The plots account for all stellar mass attributed to positive AGN feedback by considering all branches of the merger tree for each galaxy. There are notable differences between the positive AGN feedback implemented in C24 (left panel) and the new implementation in FEGA25 (right panel). First, the new positive AGN feedback in FEGA25 naturally and distinctly captures the expected redshift dependence, with higher efficiency at higher redshifts. Secondly, while in C24 there was no clear trend with stellar mass, only a larger scatter at lower mass scales (left panel), FEGA25 shows an increase in the efficiency of positive AGN feedback with stellar mass. This is particularly evident where the negative mode is also efficient, though not strong enough to heat all the cooling gas.
Figure 6: The star formation history of four sample galaxies (top 4 panels), ordered by increasing stellar mass from top to bottom, is shown from redshift $z\sim 10$ to the present time. The green line represents the secular evolution, the blue line indicates the SFR due solely to positive AGN feedback, while the red one represent the total SFR (shifted 0.1 dex upword for easier viewing). Several interesting features can be observed in these plots. From top to bottom, it is clear that the contribution of positive AGN feedback to the total SFR increases, as expected from Figure 5. This trend is evident not only in terms of the mass provided but also in the duration of activity of the positive feedback mode. Specifically, in lower-mass galaxies, the contribution is shorter and characterized by episodic bursts, whereas in higher-mass galaxies, the SFR history driven by positive AGN feedback more closely follows the total SFR, being rather continuous than broken or irregular, being longer and more sustained. In all cases, the positive feedback mode is efficient at high redshift and begins to fade as the negative feedback mode takes over the cooling process, indicating that the negative mode reaches its peak power. The two bottom panels show, instead, the distributions (median, 16th and 84th percentiles) of the SFR histories for wide samples of galaxies in the stellar mass range $11.510.5<\log M_{*}<$ 11.5 (bottom left panel), and $\log M_{*}>11.5$ (bottom right panel).
Figure 7: The fraction of mass for different components (stars, cold gas, hot gas, and ejected gas) within the virial radius, normalized by the universal baryon fraction, is shown as a function of halo mass. Different colors represent the various components, and the dashed and solid lines distinguish the predictions from FEGA24 and FEGA25, respectively. Our models are compared with the observed hot gas fraction from Popesso et al. 2024 (magenta diamonds with error bars) and a set of observed hot gas fractions from various authors (Vikhlinin et al. 2006; Arnaud et al. 2007; Sun et al. 2009; Rasmussen & Ponman 2009; Pratt et al. 2009; Mahdavi et al. 2013; Gonzalez et al. 2013; Lovisari et al. 2015; Eckert et al. 2016; Pearson et al. 2017; Eckert et al. 2019; Mulroy et al. 2019; Ragagnin et al. 2022, dark blue circles). Given the importance of sampling halos down to $\log M_{\rm{halo}}=11$, for this plot we also make use of the catalogs obtained by running the models on the merger trees extracted from YS50HR. It is important to note that, despite the large scatter in the observed data, FEGA24 fails to predict a reasonable hot gas fraction for halos smaller than large groups ($\log M_{\rm{halo}}<14$). In contrast, FEGA25, with its new hot gas ejection mode in AGN feedback, performs significantly better. This result underscores the importance of having a mechanism, such as the hot gas ejection mode, that can expel gas from the halo without significantly altering the cold gas and stellar mass content.
3.3 Positive AGN feedback and Hot Gas Ejection Modes
@@
The most evident difference between the two models lies in their trends with stellar mass. In FEGA24, the relation is flat, with a larger scatter for lower-mass galaxies and an average contribution of approximately 5%. In contrast, FEGA25 exhibits an increasing percentage and scatter with stellar mass, starting with virtually no contribution for dwarf and intermediate-mass galaxies and rising to about 5% for massive galaxies at $z=0$. At $z=3$, the percentages in FEGA24 remain consistent, albeit with larger scatter, while FEGA25 shows an increasing contribution toward lower-mass galaxies (with respect to $z=0$). This highlights a key distinction between the models: in FEGA24, redshift dependence is mildly reflected only through the scatter, whereas in FEGA25, it is a clear and systematic trend.
@@
This result is particularly significant because it aligns well with recent theoretical predictions (e.g., Silk et al. 2024), which suggest that positive AGN feedback is stronger at high redshifts—when negative feedback mechanisms are less active—and diminishes at lower redshifts, as negative feedback becomes dominant. Nevertheless, given the limited observational evidence for positive feedback (see references above), especially at high redshifts, it is important to emphasize that neither implementation can be definitively ruled out. Both positive and negative feedback modes may coexist. The advantage of the FEGA25 prescription is its ability to accurately describe the redshift dependence of positive AGN feedback.
@@
To gain an understanding of the contribution of positive AGN feedback to star formation in FEGA25, we present the SFR histories of four galaxies with increasing stellar mass, from the four top panels, in Figure 6. The green line represents the secular evolution, the blue line indicates the SFR due solely to positive AGN feedback, while the red one represent the total SFR (shifted 0.1 dex upword for easier viewing). As expected from the results in the previous figure, the overall contribution of the positive feedback mode to star formation increases with the galaxy’s stellar mass, both in terms of the mass it provides and the duration for which the positive feedback is active.
The SFR history driven by positive AGN feedback closely tracks the total SFR, displaying a rather smooth and continuous behavior rather than a broken or irregular pattern. In contrast, the total SFR is primarily sustained by secular evolution. Across all the representative cases examined, the positive feedback mode is most efficient at high redshift and gradually fades as the negative feedback mode becomes dominant, suppressing cooling more effectively. The exact redshift at which this transition occurs varies across different systems, but in general, the efficiency of the positive mode drops significantly after $z=2$, with only occasional bursts seen at later times. Notably, in several cases, the SFR triggered by positive feedback can reach—and in some instances even exceed—100 $M_{\odot}/\rm{yr}$. The bottom panels in Figure 6, instead, show the overall distributions of the SFR histories of galaxies, including the contribution of positive AGN feedback, in two different stellar mass ranges, $11.510.5<\log M_{*}<$ 11.5 (bottom left panel), and $\log M_{*}>11.5$ (bottom right panel).
@@
Finally, Figure 7 serves as the key figure of this study, clearly illustrating the role of our hot gas ejection mode. The plot shows, for each baryonic component (denoted by different colors as described in the legend), the fraction of mass relative to the total mass within the virial radius, normalized by the universal baryon fraction, $\rm{Y}_{200}$, as a function of halo mass. Dashed lines represent the predictions of FEGA24, while solid lines correspond to those of FEGA25. The hot gas fraction predicted by our models is compared with recent results from Popesso et al. (2024), shown as magenta squares, and a set of observational data from different authors (see caption), represented by dark blue circles. It is important to note that, in the latter case, the fraction was computed within the virial region using an overdensity of $\Delta=500$ (with halo masses converted to $M_{200}$ in the plot), rather than $\Delta=200$ as in our models, so these values should be considered as lower limits compared to ours.
@@
The key, and very important, distinction lies in the hot gas fraction (red lines), and by extension, in the ejected gas fraction (green lines). FEGA24 (dashed line), which lacks a hot gas ejection mode in its AGN feedback, predicts a hot gas fraction of approximately 60%-65% for halos with $\log M_{\rm{halo}}>12.2$, from small groups to large clusters. In contrast, FEGA25 (solid line), thanks to its hot gas ejection mode driven by AGN feedback, shows a considerably different hot gas fraction, more closely aligned with the observed data. The fraction of hot gas is noticeably reduced from low-mass to large groups, and while it matches the observations from Popesso et al. (2024) at the edge of the scatter in several data points, it represents a clear improvement over FEGA24.
@@
Since the cold gas and stellar components remain largely unchanged, the excess gas is directed into the ejecta (green lines). This behavior aligns with the desired outcome, as highlighted by other authors (see the discussion in Popesso et al. 2024) and as discussed in Section 1. The hot gas ejection mode in the AGN feedback has proven to be effective in reducing the hot gas within the virialized region. As emphasized in Section 1 and reiterated here, this mechanism must effectively expel gas beyond the halo’s virial radius without disrupting the internal processes that govern the cold gas and stellar components, ensuring the preservation of galaxy properties already well described by the model. That said, we observed minor differences in the amount of cold gas and stars in low-mass halos, which notably enhance the model’s ability to capture the fraction of passive dwarf galaxies.
@@
It is clear that the new achievements in our SAM stem from the introduction of a more comprehensive AGN feedback model. The idea that AGN can expel gas beyond the halo’s virial radius is almost novel in SAMs (see SHARK2, Lagos et al. 2024), though it has been quantified in numerical simulations (e.g., Pillepich et al. 2019; Angelinelli et al. 2022; Ayromlou et al. 2023; Popesso et al. 2024, among the most recent studies). Popesso et al. (2024) provides an updated and detailed analysis of this phenomenon, comparing numerical simulation results with observational data (e.g., eROSITA eFEDS field data, Brunner et al. 2022). As discussed in Section 1, most simulations struggle to reproduce the observed fraction of hot gas within the virialized region over a broad range of halo masses, particularly at group scales. This shortfall underscores the need for an additional mechanism capable of pushing hot gas beyond the virial radius.
@@
However, further investigation is needed for a particular reason. While FEGA25 performs considerably better than its predecessor, its predictions still tend toward the upper limits of the observed range, suggesting that an even more efficient mechanism may be required. As described in Section 2.1.3, we can assume that the residual energy, after reheating the cooling gas in the negative feedback mode, is sufficient to expel some of the hot gas beyond the virial radius. While this is a plausible and valuable approach, it represents an extreme scenario that warrants a more in-depth investigation. We plan to explore this topic further in a forthcoming paper (Contini et al., in prep.), which will focus on the evolution of the baryon content in halos and its dependence on cosmic time.